OpenStudy (anonymous):

one fourth of a number q is 3 more than 3 times a number w solve for q . help please

jimthompson5910 (jim_thompson5910):

"one fourth of a number q is 3 more than 3 times a number w" translates to q/4 = 3w+3 Now let's solve for q q/4 = 3w+3 q = 4(3w+3) q = 12w+12

OpenStudy (whynot):

q/4\[q/4=3w+3\]

OpenStudy (phi):

one fourth of a number q is 3 more than 3 times a number w. one fourth means 1/4 1/4 of a number q means (1/4) q or q/4 another example: 1/2 of 2 --> (1/2)*2 or 2/2 or 1 q/4 is means q/4 = is --> equal, or = Example: 2*2 is 4 --> 2*2 = 4 sometimes "as much as" means equal q/4 = 3 more than 3 more than means 3+ Example: 3 more than 2 --> 3+2 q/4 = 3 + 3 times a number w 3 times w means 3 w q/4 = 3 + 3 w solve for q . now it is algebra.